# On zeros of polynomials associated with Heun class equations

### v2

@@ 7. Concluding remarks
In [14], a perturbative approach to the Calogero-Moser-Sutherland system, which is related with Heun’s differential equation, was performed, and it was applied to the Lamé equation in [16]. It would be interesting to compare the results in [14, 16] to the ones in this paper. It is known that the Lamé equation has the finite-gap property (see [17] and references therein). The numerical calculations in Section 5 would produce approximately a part of the spectral of the Lamé equation. Details on this aspect should be revealed in a near future. Some examples of the numerical calculations for the Mathieu equation and for the Whittaker-Hill equation were discussed, and the results would be approximately a part of the spectral of those equations. We would be able to carry out numerical calculations for the Heun equation, the singly confluent Heun equation and the reduced singly confluent Heun equation with the condition $\gamma\neq 1/2$ or $\delta\neq 1/2$, and we expect further studies for the spectral of them.

### v3

@@ 7. Concluding remarks
In [14], a perturbative approach to the Calogero-Moser-Sutherland system, which is related with Heun’s differential equation, was performed, and it was applied to the Lamé equation in [16]. It would be interesting to compare the results in [14, 16] to the ones in this paper. It is known that the Lamé equation has the finite-gap property (see [17] and references therein). The numerical calculations in Section 5 would produce a part of the spectral of the Lamé equation as an approximation. Details on this aspect should be revealed in a near future. Some examples of the numerical calculations for the Mathieu equation and for the Whittaker-Hill equation were discussed in Section 6, and the results would produce a part of the spectral of those equations as an approximation. It would also be possible to carry out numerical calculations for the Heun equation, the singly confluent Heun equation and the reduced singly confluent Heun equation with the condition $\gamma\neq 1/2$ or $\delta\neq 1/2$, and we expect further studies for the spectral of them.
Although we obtained exact results on the zeros of the polynomial $c_{m+1}(B)$ from the perturbative approach in Section 4, the zeros were discussed numerically in Sections 5 and 6, and we did not have exact estimates for error bounds. In other words, we discussed that some zeros of $c_{m+1}(B)$ are sufficiently close to some zeros of $c_{m^{\prime}+1}(B)$ numerically for some values $m$ and $m^{\prime}$ by examples, but we did not discuss exact estimates of the zeros of $d_{2}(B)$ and limits of some zeros of $c_{m+1}(B)$ as $m\to\infty$. Note that the zeros of $d_{2}(B)$ are directly related to the global structure of the solutions to the corresponding Heun-type equations. To obtain exact estimates, it would be necessary to develop the analytic argument of Schäfke and Schmidt [11] to be fit with the uniform convergence on compact sets with respect to the parameter $B$ for Heun-class equations. The approach by Volkmer [18] might be useful to discuss convergence of the zeros.